Using samples of 500credit card statements, an auditor found the following: Use Table-A. Sample 1 2 3 4 Number with errors 5 3 6 9 a. What is the 3 sigma control limit for this process
Answer #1
Number of errors Error proportion
5 5/500 = 0.01
3 0.006
6 0.06
9 0.09

Mean of errors = Sum of error propotions / 4 = 0.0415

Std deviation of proportions = sqrt { Mean of proportion * (1- Mean of proportion}/n } = sqrt{ (0.0415) * (1-0.0415) / 500 } = 0.0087

Three sigma control process = Mean + / - Z*Std deviation

Z= 3

UCL = 0.0415 + 3*0.0087 = 0.676

LCL = 0.0415 - 3*0.0087 = 0.0154

P (1-P